OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y =x2 - 2x - 48. A. -49 B. 49 C. 48 D. -48

OpenStudy (abb0t):

To find the minimum or maxium, you need to take the first derivative of the given function.

OpenStudy (anonymous):

i dont get it

OpenStudy (abb0t):

Do you know how to take derivatives?

OpenStudy (jdoe0001):

@miryluslinares are you doing calc?

OpenStudy (anonymous):

no i am doing algebra 2

OpenStudy (abb0t):

If this is not for calculus, you need to put it in standard form: y=A(x-h)\(^2\)+k, where (h,k)=(x,y) are the coordinates of the vertex

OpenStudy (abb0t):

That means, completing the square.

OpenStudy (jdoe0001):

I assume you've done parabolas, so, the minimum or maximum, will be at the LOWEST or HIGHEST point of the parabola, and that occurs at its vertex you can find the vertex of a parabola \(\bf ax^2+bx+c\) at \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (jdoe0001):

or you can also as abb0t said, complete the square and get the "vertex form" of the parabola, that'd work too

OpenStudy (jdoe0001):

keep in mind that, what they mean by "minimum" or "maximum", is just the "hump" the parabola has, which is also the "vertex"

OpenStudy (anonymous):

how would i plug it in that that equation its because i am taking an online class and i am lost

OpenStudy (jdoe0001):

\(\large \begin{array}{lllllll} &a&&b&&c\\ y =x^2 - 2x - 48 \implies y =&1&x^2 &- 2&x &- 48 \end{array}\\\quad \\ \bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\) notice the coefficients

OpenStudy (anonymous):

ohhh ok i get it thank you very much :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

@jdoe0001 so its D right?

OpenStudy (jdoe0001):

well, they're asking for the value for "y", that is, f(x) at the vertex, so, they want the y-coordinate of it \(\bf \large \qquad x \qquad \qquad y\\ \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (jdoe0001):

so, what did you get for "y"?

OpenStudy (anonymous):

i got -48.5

OpenStudy (anonymous):

am i doing it right?

OpenStudy (jdoe0001):

let's see \(\large \begin{array}{lllllll} &a&&b&&c\\ y =x^2 - 2x - 48 \implies y =&1&x^2 &- 2&x &- 48 \end{array}\\\quad \\ \bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\\ \quad \\ c-\cfrac{b^2}{4a} \implies (-48)-\cfrac{(-2)^2}{4(1)}\)

OpenStudy (anonymous):

nevermind i got -49 thank you for taking the time to help me

OpenStudy (jdoe0001):

yw